OpenStudy (anonymous):

solve for x round to nerest 2 decimal places ln=-2.5 and 2x log base3 x+4=1

OpenStudy (anonymous):

x+4= 3^(1/2)

OpenStudy (anonymous):

x=sqrt(3)-4

OpenStudy (anonymous):

-2.27

OpenStudy (anonymous):

thank you and what about ln=-2.5

OpenStudy (anonymous):

there to diffrent problems

OpenStudy (anonymous):

ln=-2.5? where is the x?

OpenStudy (anonymous):

x=e^-2.5=0.08